OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 (3, 2) (2, 3) (-2, -3) (-2, 3)

OpenStudy (anonymous):

well uh, 4x+y = 11 so y = 11-4x. use y = 11 - 4x in x+2y = 8 x + 2(11-4x) = 8. Solve accordingly for x, y

jimthompson5910 (jim_thompson5910):

solve the first equation for y 4x + y = 11 y = -4x + 11 ---------------- then plug this into the second equation x + 2y = 8 x + 2(-4x + 11) = 8 x - 8x + 22 = 8 I'll let you finish up